OpenStudy (anonymous):

find all of the real or imaginary solutions of each equation by factoring 6x^3-2x^2+4x=0

OpenStudy (anonymous):

6x^3-2x^2+4x=0 x(6x^2-2x+4) =0 now use the factoring formula (-b+- sqrt b^2 -4ac all over 2a) You get that the roots are x=0 x= (1/4)+2i(sqrt 23) x= (1/4)-2i(sqrt 23)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yup. Medal?